OpenStudy (anonymous):

Fruit flies are placed in a half-pint milk bottle with a banana (for food) and yeast plants (for food and to provide a stimulus to lay eggs). Suppose that the fruit fly population after t days is given by:

OpenStudy (anonymous):

\[p(t)=\frac{ 230 }{ 1+56.5e ^{-0.37t} }\] What is the carrying capacity of the half pint bottle? (i.e, What is p(t) as t approaches ∞?)

OpenStudy (anonymous):

those question marks represent infinity. i don't know why it didn't show

OpenStudy (perl):

what does e^-infinity represent

OpenStudy (anonymous):

what do you mean?

OpenStudy (ribhu):

230

OpenStudy (anonymous):

can you briefly explain?

OpenStudy (ribhu):

P(t) would be 230 as t approaches infinite

OpenStudy (anonymous):

because my teacher told me we have to write like a little paragraph on explaining it.

OpenStudy (ribhu):

yeah i can, see when t tends to infinity the denominator becomes 1 as, 56.5e^-0.37t would be zero

OpenStudy (anonymous):

wait so how does that prove that the carrying capacity is 230?

OpenStudy (ribhu):

the same thing is asked in the question? the value of P at t=infinite.

OpenStudy (perl):

right

OpenStudy (anonymous):

I'm sorry I'm confused...

OpenStudy (perl):

you want lim(t->oo) 230 / ( 1 + 56.5 * e^ (-.37*t)

OpenStudy (ribhu):

why see at t=0 the value is 4

OpenStudy (anonymous):

so 4 is the initial amount of fruit flies? how does that help

OpenStudy (ribhu):

and the fruit flies will occupy the whole bottle. they will not grow when the no banana is left

OpenStudy (ribhu):

there is banana in the half pint bottle and fruit flies grow with the function given. they will grow till they have eaten whole banana.

OpenStudy (perl):

lim(t->oo) 230 / ( 1 + 56.5 * e^ (-.37*t) = 230 / ( 1 + 56.5 * e^(-.37 * oo) = 230 / ( 1 + 56.5 *0) =230/(1 + 0) =230

OpenStudy (ribhu):

assumption that at t=infinity they must have consumed whole banana and ability to multiply again will be finished

OpenStudy (ribhu):

this would be my set of explanation for this problem.

OpenStudy (perl):

the question is not asking about the banana, it is asking the total population of fruit flies

OpenStudy (perl):

but youre right, they would probably eat the banana and the yeast

OpenStudy (perl):

:)